OpenStudy (anonymous):

Thoroughbred Bus Company finds that its monthly costs for one particular year were given by C(t) = 10,000 + t2 dollars after t months. After t months the company had P(t) = 1,000 + t2 passengers per month. How fast is its cost per passenger changing after 8 months? HINT [See Example 8(b).] (Round your answer to two decimal places.) $_____________ per month

OpenStudy (anonymous):

so plug 8 into both equations

OpenStudy (anonymous):

i did but it's a negative #

OpenStudy (anonymous):

C(t)/P(t)

OpenStudy (anonymous):

i think i should also take the derivative of that

OpenStudy (anonymous):

depends what does 8(b) look like?

OpenStudy (anonymous):

the question ask "how fast" so it's asking the change of rate

OpenStudy (anonymous):

so yes you would

OpenStudy (anonymous):

ok. thank you, though.

OpenStudy (anonymous):

you got it from here

OpenStudy (anonymous):

nope, the answer was not right

OpenStudy (anonymous):

here is a similar problem http://answers.yahoo.com/question/index?qid=20111107225609AARAMOb

OpenStudy (anonymous):

Haha, i have looked that too. but it wasn't right

OpenStudy (anonymous):

-2250/17689

OpenStudy (anonymous):

that was the same as mine -0.127

OpenStudy (anonymous):

hmmm dont know what to tell ya

OpenStudy (anonymous):

@hartnn he has all the answers

OpenStudy (anonymous):

really? can i ask @hartnn, then?

OpenStudy (anonymous):

ya he should be here i a few he is helping someone else

OpenStudy (anonymous):

he don't know how to do it

OpenStudy (anonymous):

did you ask?

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

lets try @Jonask he might be able to

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ C }{ month }=10000+t^2\] \[\frac{ P }{ month }=1000+t^2\] \[\frac{ C }{ P }=?\] \[\frac{ C }{ month } \div \frac{ P }{ month }=\frac{ C }{ P }\]

OpenStudy (anonymous):

\[\frac{ C }{ P }=(10000+t^2)\div(1000+t^2)\] t=8

OpenStudy (anonymous):

i am thinking about take the derivative of C(t)/P(t)

OpenStudy (anonymous):

what i got is a negative #

OpenStudy (anonymous):

\[\frac{ dC }{ dt }\div\frac{ dP }{ dt}=\frac{ dC }{ dP }\]

OpenStudy (anonymous):

yap, the answer is -0.13, but it's wrong

OpenStudy (anonymous):

\[\frac{ dC }{ dP }=(10000+t^2)(1000+t^2)\] \[\frac{ 10000+8^2 }{ 1000+8^2 }=\frac{ 10064 }{ 1064 }\]

OpenStudy (anonymous):

\[\frac{ dC }{ dP }=(10000+t^2) \div (1000+t^2)\]

OpenStudy (anonymous):

which is also wrong

OpenStudy (anonymous):

so\[\frac{ dC }{ dP }=\frac{ 10000+t^2 }{ 1000+t^2 }\]

OpenStudy (anonymous):

whats your answer

OpenStudy (anonymous):

-0.13

OpenStudy (anonymous):

1258/133

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

c(t) has to be derived

OpenStudy (anonymous):

yap, that's how i got the negative #

OpenStudy (anonymous):

2t but remember cost is naturally negetive

OpenStudy (anonymous):

but what i got is wrong

OpenStudy (anonymous):

ok guys good luck hope you get it but i g2g to bed

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[C(t)=\frac{ 10000+t^2 }{ 12 }\] per month

OpenStudy (anonymous):

wish i could have helped more, but atleast i found someone to help so we will call it even. thanks jonask

OpenStudy (anonymous):

but the question doesn't ask for each month

OpenStudy (anonymous):

the function given is for one year,so for a month it will be /12

OpenStudy (anonymous):

didnt see that good catch

OpenStudy (anonymous):

\[\frac{ dC }{ dP }=\frac{ 2(8) }{ 1000+64 }\] anyway how is this -

OpenStudy (anonymous):

do you know the answer @laulungyuen

OpenStudy (anonymous):

why is the ansewr in the form $...month instead of $...per passenger

OpenStudy (anonymous):

i think i should use quotient rule

OpenStudy (anonymous):

\[(\frac{ 10000+64 }{ 12 })\div(1000+64)=\frac{ 10064 }{ 12(1064) }\]

OpenStudy (anonymous):

0,788

OpenStudy (anonymous):

but the equation you gave me is not related to the derivative, though

OpenStudy (anonymous):

we have to derive this\[\frac{ dC }{ dP }\]because question says "How fast " \[\frac{ dC }{ dP }=\frac{ 10000+t^2 }{ 1000+t^2 }\] \[\frac{ d(\frac{ dC }{ dP } )}{ dt }=\frac{ 2t(1000+t^2)-2t(10000+t^2) }{ (1000+t^2) }\]

OpenStudy (anonymous):

0,127

OpenStudy (anonymous):

-0,127

OpenStudy (anonymous):

yap, but it's wrong

OpenStudy (anonymous):

how much is the answer

OpenStudy (anonymous):

do you have 8(b)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

how do you know its wrong if you dont have the answer

OpenStudy (anonymous):

im doing online assignment, so if the answer is wrong it will tell me

OpenStudy (anonymous):

try -0,01059 i have to go to school ,,,later

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

morning class?

OpenStudy (anonymous):

yes,its 9am

OpenStudy (anonymous):

oh ok thanks